# Orthogonal projections of hypercubes

### v2

@@ IV.4 Brief review and preview
When two PC loadings cannot capture all the weighted vertices, the missed weighted vertices can be placed near the origin of the projected space. This centrality of the projected vertices is discussed further in Sec. VI.2 and its effect on the inner-product error is discussed in Sec. VI.3.
@@ VI.2 Centrality of the projection
We first consider the number of vertices along the Hamming-distance axis of the Hamming projection. The projected coordinates with Hamming distance $D_{\mathrm{H}}=D_{\mathrm{H}}\left(-\bm{\xi},\bm{s}\right)$ are given by Eq. (18), $r\left(D_{\mathrm{H}}\right)\in\left\{-\sqrt{N}+\frac{2}{\sqrt{N}}D_{\mathrm{H}}\right\}_{D_{\mathrm{H}}=0}^{N}$. Because the Hamming distance is the number of unmatched elements in the binary vectors, the binomial coefficient $\binom{N}{D_{\mathrm{H}}}\coloneqq\frac{N!}{D_{\mathrm{H}}!\left(N-D_{\mathrm{H}}\right)!}$ gives the number of vertices at each possible $r\left(D_{\mathrm{H}}\right)$. The normalized number of vertices with the Hamming distance $D_{\mathrm{H}}$ is $\varrho\left(D_{\mathrm{H}}\right)=\binom{N}{D_{\mathrm{H}}}\frac{1}{2^{N}}=\binom{N}{D_{\mathrm{H}}}\left(\frac{1}{2}\right)^{D_{\mathrm{H}}}\left(1-\frac{1}{2}\right)^{N-D_{\mathrm{H}}}$, i.e., the binomial distribution. With large $N$, $\varrho\left(D_{\mathrm{H}}\right)$ is asymptotically a Gaussian distribution with mean $\frac{N}{2}$ and variance $\frac{N}{4}$,
$\varrho\left(D_{\mathrm{H}}\right)\sim\frac{1}{\sqrt{2\uppi\frac{N}{4}}}\exp\left[-\frac{1}{2}\frac{\left(D_{\mathrm{H}}-\frac{N}{2}\right)^{2}}{\frac{N}{4}}\right].$ (23)
@@ VI.3 Inner-product error bounds of projections
Notice that the corners of Fig. 8(b) have a slightly higher bound, but here we show that such regions cannot be used as projected coordinates. The possible range of $r_{1}\left(\bm{s}\right)$ and $r_{2}\left(\bm{s}\right)$ is $\left[-\sqrt{N},\sqrt{N}\right]$. In Appendix D, however, we show that projected coordinates are limited to a specific region of
@@
What we show here is the reason that many vertices are projected around the origin in the orthogonal projection—whether by PCA or not—of hypercubes. Upon linear projection, hypercubic vertices concentrate near the origin of the projected space in an approximate Gaussian shape. This is why the inner-product error gets a dominant contribution from the vertices located around the center of the projected space. Indeed, the upper bound of the inner-product error is the largest when the vertices are projected around the origin. Our discussion in Secs. VI.2 and VI.3 entirely relies on the assumption that the hypercubic vertices are projected to form a Hamming projection, which is a rough approximation and might not be the case in general. The full properties of such non-Hamming projections cannot be revealed analytically, but as we show in Sec. V.2, the qualitative aspects of such projections are expected to be similar to those of the Hamming projection. Thus, the quality of the projection of hypercubes is typically worse around the center of the projected space.
@@ VII.2 Projecting the hypercubic energy landscape of the kagome spin-ice system
Figure 10: Hypercubic energy landscape of finite artificial spin-ice systems and interaction networks colored by the angle of the biplot basis. (a) The hypercubic energy landscape of one-ring (a1), two-ring (a2), and three-ring (a3) systems by PC1 and PC2. For clarity of the figure, only the states with the lowest and the second-lowest energy are shown. The energy of the state colors vertices (blue for the lowest energy and red for the second-lowest energy). Biplot bases are shown on the top right, and their color indicates the angle of the biplot basis. Notice that in (a2) and (a3), several biplot bases are overlapped. One example pathway from a ground state to another ground state is shown as colored hypercubic edges in (a1). The colors of these lines are the same as the color of the corresponding biplot basis. Examples of states are shown as insets in (a1), (a2), and (a3). See the caption of Fig. 9 for the definition of state visualization on the interaction network. (b) PC1 loading of one-ring (b1), two-ring (b2), and three-ring (b3) systems. The color of the bar matches the color of the corresponding biplot basis in (a). (c) Same as (b) but for PC2. (d) The Ising spin interaction networks of one-ring (d1), two-ring (d2), and three-ring (d3) systems, where the nodes are colored by the angle of the corresponding biplot basis in (a).
@@ VII.3 The hypercubic energy landscape of the one-ring system
We first examine the one-ring system of Figs. 9(a1) and 9(b1). In Fig. 10(a1), we show the possible pathway from a ground state $\begin{bmatrix}\circ&\bullet&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ to the globally spin-flipped state $\begin{bmatrix}\bullet&\circ&\circ&\bullet&\bullet&\circ\end{bmatrix}^{\top}$ by connecting the vertices with the lowest energy and the second-lowest energy using hypercubic edges. (See Figs. 9(b1) or 10(d1) for labeling of spins.) In this example, the hypercubic edge corresponds to the state transition with a single spin flip. The biplot bases are shown at the top right of Fig. 10(a1), and its color indicates the angle. We emphasize one particular pathway between two ground states by coloring the edges following the biplot bases in Fig. 10(b1). We found the state transition pathway is collective: the flipping order is restricted to keep the energy low. In a one-ring system, there are $6!=720$ possible pathways from one ground state to another ground state, but the number of most probable pathways is limited, as shown in Fig. 10(a1), and it is reduced to $966\times 2^{4}\times 1=$ 96 [69]. This reduction arises from the interaction network of the system and the associated hypercubic energy landscape.
@@
To interpret the relation between PC loading and the interaction network, we visualize the loading of PC1 and PC2 in Figs. 10(b1) and 10(c1), respectively, colored by the angle of the biplot bases in Fig. 10(a1). We confirmed that the element-wise sign of PC1 corresponds to the ground state [Fig. 10(b1)]. The element-wise sign of PC2 loading corresponds to the state perpendicular to the ground state, which has a Hamming distance of $\frac{N}{2}$ from the ground state [Fig. 10(c1)].
@@ VII.4 The hypercubic energy landscape of the two-ring system
The second example is the transition pathway from the ground state on the upper left side (around $\begin{bmatrix}-3&0.3\end{bmatrix}^{\top}$), to the ground state on the top (around $\begin{bmatrix}0&3.2\end{bmatrix}^{\top}$), i.e., state transition from $\begin{bmatrix}\circ&\bullet&\circ&\bullet&\bullet&\bullet&\circ&\circ&\bullet&\circ&\bullet\end{bmatrix}^{\top}$ to $\begin{bmatrix}\bullet&\circ&\circ&\bullet&\circ&\bullet&\circ&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$, which requires the flipping of spins 1, 2, 5, 8, and 9 [see biplot bases of Fig. 10(a2) and color of Figs. 10(b2), 10(c2), and 10(d2)]. This projection of the hypercube indicates that state transition happens with the correlated spin flips on the left-half part of the system (spins 1, 2, 5, 8, and 9). Among the left-half part of the system, spins 2 and 9 are more likely to flip at the beginning of the pathway. This difference arises from the unsatisfied interactions: $J_{3,9}$ and $J_{6,2}$. Spins 2 and 9 receive weaker local fields than spins 1, 5, and 8 because they have interactions involving spins 3, 6, and 10. Thus, spins 2 and 9 are expected to flip more than spins 1, 5, and 8. Because of the frustrated interactions involving spins 3, 6, and 10, spins 2 or 9 are the most probable to flip first. Then, the flips happen in the order of spins 2, 1, 5, 8, and 9 if spin 2 flips first, or in the order of spins 9, 8, 5, 1, and 2 if spin 9 flips first. Similar to the one-ring system in Fig. 10(a1), the number of most probable pathways is limited, but the constraints are stricter and fewer pathways are possible.
@@
To deepen our understanding of PC loading and interaction networks, we show the loading of PC1 and PC2 in Figs. 10(b2) and 10(c2), coloring them by the angle of the biplot bases in Fig. 10(a2). Again, we confirmed that the sign of PC1 loading [Fig. 10(b2)] corresponds to the ground state with the largest magnitude of the PC1 score, but spin 6 is not determined. Thus, the frustrated dynamics of spin 6 are captured by PC1. The element-wise sign of PC2 loading [Fig. 10(c2)] corresponds to the ground state with the largest magnitude of the PC2 score. Notice that the sixth element of PC2 loading has a slightly higher magnitude than the others, indicating the importance of spin 6. If spin 6 flips when the system is in the ground state $\operatorname{sgn}\left(\bm{u}_{2}\right)$, the energy of the system increases more than when another spin flips.
@@
The interaction network with nodes colored by the angle of the biplot bases in Fig. 10(d2) shows the inter-spin correlation captured by the biplot bases. The biplot bases capture the (anti)correlation of spins 1, 2, 5, 8, and 9 [yellow and blue arrows in Fig. 10(a2) top right] which corresponds to the left ring of the interaction network in Fig. 10(d2). The biplot bases also capturs the (anti)correlation of spins 3, 4, 7, 10, and 11 [green and purple arrows in Fig. 10(a2) top right] which corresponds to the right ring of the interaction network in Fig. 10(d2). The hypercubic energy landscape using PCA provides insight into the dynamics through the correlation of spins.
@@ VII.5 The hypercubic energy landscape of the three-ring system
Unlike the one-ring and two-ring systems, the structure of pathways is more complex, but we can still infer the state transition from the projection. For example, the pathway from the ground state $\begin{bmatrix}\circ&\bullet&\circ&\bullet&\bullet&\circ&\circ&\circ&\circ&\bullet&\bullet&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ (around $\begin{bmatrix}-3&0.9\end{bmatrix}^{\top}$) to $\begin{bmatrix}\circ&\bullet&\bullet&\circ&\bullet&\circ&\bullet&\circ&\circ&\bullet&\circ&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ (around $\begin{bmatrix}-1.5&3.2\end{bmatrix}^{\top}$) is a transition with the correlated spin flips of spins located in the lower right hexagon (spins 3, 4, 7, and 11). Pathways parallel to this pathway involve the same spin flips. From the biplot bases on the top right of Fig. 10(a3), one can generally know that by following the angle of each biplot basis, the transitions involve spins on the top hexagon (spins 12, 13, 14, and 15) being flipped: the pathways are lines from the left top to the right bottom, which are parallel to the biplot bases of spins 12, 13, 14, and 15. Likewise, horizontal pathways are used when spins in the lower left hexagon (spins 1, 2, 5, and 8) are flipped, and vertical pathways are followed when spins in the inverted triangle at the center (spins 6, 9, and 10) are flipped. Similar to the one-ring and two-ring systems, the pathway is a collective flip of spins reflecting the interaction network. Notice that this projection of the hypercubic energy landscape [Fig. 10(a3)] is a partial Hamming projection: the PC1 score is equivalent to the Hamming distance from the ground state on the left side (around $\begin{bmatrix}-3&0\end{bmatrix}^{\top}$) or the right side (around $\begin{bmatrix}3&0\end{bmatrix}^{\top}$) but without considering spins 6, 9, and 10.
@@
To understand the relation between PC loading and the interaction network, in particular, the relation between PC loading and the frustration of the system, we color the loading of PC1 and PC2 in Figs. 10(b3) and 10(c3), following the angle of the biplot bases in Fig. 10(a3). We confirmed that the element-wise sign of PC1 [Fig. 10(b3)] corresponds to the ground state group containing six ground states, but the states of spins 6, 9, and 10 are not determined. These spins involve a frustrated cycle of the interaction network. PC2 loading in Fig. 10(c3) shows that spins 1, 2, 5, and 8 do not contribute, and those spins belong to the lower left ring. Moreover, PC2 loading of spins 6, 9, and 10 has slightly higher or lower magnitudes than the others, indicating the uncommon contribution of spins 6, 9, and 10, suggesting the frustration-related correlation.
@@
These insights are consistent with the interaction network with colored nodes in Fig. 10(d3): the PC loading captures the correlation arising from the interaction network. Nevertheless, spins 6 and 9 have exactly the same color: their biplot bases are the same, but it can be another combination of spins such as 6 and 10. This breaks the symmetry of the system—PC2 ignores the lower left hexagon of the interaction network even though the system has the same other structure—and we examine how the PC loading reflects the symmetry of the system.

### v3

@@ IV.4 Brief review and preview
When two PC loadings cannot capture all the weighted vertices, the missed weighted vertices can be placed near the origin of the projected space. This centrality, often seen in studies of Ising spin system [111, 112, 113], of the projected vertices is discussed further in Sec. VI.2 and its effect on the inner-product error is discussed in Sec. VI.3.
@@ VI.2 Centrality of the projection
We first consider the number of vertices along the Hamming-distance axis of the Hamming projection. The projected coordinates with Hamming distance $D_{\mathrm{H}}=D_{\mathrm{H}}\left(-\bm{\xi},\bm{s}\right)$ are given by Eq. (18), $r\left(D_{\mathrm{H}}\right)\in\left\{-\sqrt{N}+\frac{2}{\sqrt{N}}D_{\mathrm{H}}\right\}_{D_{\mathrm{H}}=0}^{N}$. Because the Hamming distance is the number of unmatched elements in the binary vectors, the binomial coefficient $\binom{N}{D_{\mathrm{H}}}\coloneqq\frac{N!}{D_{\mathrm{H}}!\left(N-D_{\mathrm{H}}\right)!}$ gives the number of vertices at each possible $r\left(D_{\mathrm{H}}\right)$. The normalized number of vertices with the Hamming distance $D_{\mathrm{H}}$ is $\varrho\left(D_{\mathrm{H}}\right)=\binom{N}{D_{\mathrm{H}}}\frac{1}{2^{N}}=\binom{N}{D_{\mathrm{H}}}\left(\frac{1}{2}\right)^{D_{\mathrm{H}}}\left(1-\frac{1}{2}\right)^{N-D_{\mathrm{H}}}$, i.e., the binomial distribution. With large $N$, de Moivre–Laplace theorem states that $\varrho\left(D_{\mathrm{H}}\right)$ is asymptotically a Gaussian distribution with mean $\frac{N}{2}$ and variance $\frac{N}{4}$,
$\varrho\left(D_{\mathrm{H}}\right)\simeq\frac{1}{\sqrt{2\uppi\frac{N}{4}}}\exp\left[-\frac{1}{2}\frac{\left(D_{\mathrm{H}}-\frac{N}{2}\right)^{2}}{\frac{N}{4}}\right].$ (23)
@@
What we show here is the reason that many vertices are projected around the origin in the orthogonal projection—whether by PCA or not—of hypercubes. Upon linear projection, hypercubic vertices concentrate near the origin of the projected space roughly following Gaussian shape.
@@ VI.3 Inner-product error bounds of projections
Notice that the corners of Fig. 8(b) have a slightly higher bound, but here we show that such regions cannot be used as projected coordinates. Assume we have a Hamming projection with two weighted states. From Eq. (14), the possible range of a single-dimensional projected coordinate is $-\sqrt{N}\leq r_{i}\left(\bm{s}\right)\leq\sqrt{N}$. In Appendix D, however, we show that a two-dimensional projected coordinate is limited to a specific region of
@@
This Sec. VI.3 explains why the inner-product error gets a dominant contribution from the vertices located around the center of the projected space. Indeed, the upper bound of the inner-product error is the largest when the vertices are projected around the origin. This tendency is consistent with the numerical results in the previous work [97]. Together with the results in Sec. VI.2, we conclude that the vertices projected around the center of the projected space contribute most to the inner-product error.
Our discussion in Secs. VI.2 and VI.3 entirely relies on the assumption that the hypercubic vertices are projected to form a Hamming projection, which is a rough approximation and might not be the case in general. The full properties of such non-Hamming projections cannot be revealed analytically, but as we show in Sec. V.2, the qualitative aspects of such projections are expected to be similar to those of the Hamming projection. Thus, the quality of the projection of hypercubes is typically worse around the center of the projected space.
@@ VII.2 Projecting the hypercubic energy landscape of the kagome spin-ice system
Figure 10: Hypercubic energy landscape of finite artificial spin-ice systems and interaction networks colored by the angle of the biplot basis. (a) The hypercubic energy landscape of one-ring (a1), two-ring (a2), and three-ring (a3) systems by PC1 and PC2. For clarity of the figure, only the states with the lowest and the second-lowest energy are shown. The energy of the state colors vertices (blue for the lowest energy and red for the second-lowest energy). Biplot basis are shown on the top right, and their color indicates the angle of the biplot basis. Notice that in (a2) and (a3), several biplot basis are overlapped. One example pathway from a ground state to another ground state is shown as colored hypercubic edges in (a1). The colors of these lines are the same as the color of the corresponding biplot basis. Examples of states are shown as insets in (a1), (a2), and (a3). See the caption of Fig. 9 for the definition of state visualization on the interaction network. (b) PC1 loading of one-ring (b1), two-ring (b2), and three-ring (b3) systems. The color of the bar matches the color of the corresponding biplot basis in (a). (c) Same as (b) but for PC2. (d) The Ising spin interaction networks of one-ring (d1), two-ring (d2), and three-ring (d3) systems, where the nodes are colored by the angle of the corresponding biplot basis in (a).
@@ VII.3 The hypercubic energy landscape of the one-ring system
We first examine the one-ring system of Figs. 9(a1) and 9(b1). In Fig. 10(a1), we show the possible pathway from a ground state $\begin{bmatrix}\circ&\bullet&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ to the globally spin-flipped state $\begin{bmatrix}\bullet&\circ&\circ&\bullet&\bullet&\circ\end{bmatrix}^{\top}$ by connecting the vertices with the lowest energy and the second-lowest energy using hypercubic edges. (See Figs. 9(b1) or 10(d1) for labeling of spins.) In this example, the hypercubic edge corresponds to the state transition with a single spin flip. The biplot basis are shown at the top right of Fig. 10(a1), and its color indicates the angle. We emphasize one particular pathway between two ground states by coloring the edges following the biplot basis in Fig. 10(b1). We found the state transition pathway is collective: the flipping order is restricted to keep the energy low. In a one-ring system, there are $6!=720$ possible pathways from one ground state to another ground state, but the number of most probable pathways is limited, as shown in Fig. 10(a1), and it is reduced to $966\times 2^{4}\times 1=$ 96 [70]. This reduction arises from the interaction network of the system and the associated hypercubic energy landscape.
@@
To interpret the relation between PC loading and the interaction network, we visualize the loading of PC1 and PC2 in Figs. 10(b1) and 10(c1), respectively, colored by the angle of the biplot basis in Fig. 10(a1). We confirmed that the element-wise sign of PC1 corresponds to the ground state [Fig. 10(b1)]. The element-wise sign of PC2 loading corresponds to the state perpendicular to the ground state, which has a Hamming distance of $\frac{N}{2}$ from the ground state [Fig. 10(c1)].
@@ VII.4 The hypercubic energy landscape of the two-ring system
The second example is the transition pathway from the ground state on the upper left side (around $\begin{bmatrix}-3&0.3\end{bmatrix}^{\top}$), to the ground state on the top (around $\begin{bmatrix}0&3.2\end{bmatrix}^{\top}$), i.e., state transition from $\begin{bmatrix}\circ&\bullet&\circ&\bullet&\bullet&\bullet&\circ&\circ&\bullet&\circ&\bullet\end{bmatrix}^{\top}$ to $\begin{bmatrix}\bullet&\circ&\circ&\bullet&\circ&\bullet&\circ&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$, which requires the flipping of spins 1, 2, 5, 8, and 9 [see biplot basis of Fig. 10(a2) and color of Figs. 10(b2), 10(c2), and 10(d2)]. This projection of the hypercube indicates that state transition happens with the correlated spin flips on the left-half part of the system (spins 1, 2, 5, 8, and 9). Among the left-half part of the system, spins 2 and 9 are more likely to flip at the beginning of the pathway. This difference arises from the unsatisfied interactions: $J_{3,9}$ and $J_{6,2}$. Spins 2 and 9 receive weaker local fields than spins 1, 5, and 8 because they have interactions involving spins 3, 6, and 10. Thus, spins 2 and 9 are expected to flip more than spins 1, 5, and 8. Because of the frustrated interactions involving spins 3, 6, and 10, spins 2 or 9 are the most probable to flip first. Then, the flips happen in the order of spins 2, 1, 5, 8, and 9 if spin 2 flips first, or in the order of spins 9, 8, 5, 1, and 2 if spin 9 flips first. Similar to the one-ring system in Fig. 10(a1), the number of most probable pathways is limited, but the constraints are stricter and fewer pathways are possible.
@@
To deepen our understanding of PC loading and interaction networks, we show the loading of PC1 and PC2 in Figs. 10(b2) and 10(c2), coloring them by the angle of the biplot basis in Fig. 10(a2). Again, we confirmed that the sign of PC1 loading [Fig. 10(b2)] corresponds to the ground state with the largest magnitude of the PC1 score, but spin 6 is not determined. Thus, the frustrated dynamics of spin 6 are captured by PC1. The element-wise sign of PC2 loading [Fig. 10(c2)] corresponds to the ground state with the largest magnitude of the PC2 score. Notice that the sixth element of PC2 loading has a slightly higher magnitude than the others, indicating the importance of spin 6. If spin 6 flips when the system is in the ground state $\operatorname{sgn}\left(\bm{u}_{2}\right)$, the energy of the system increases more than when another spin flips.
@@
The interaction network with nodes colored by the angle of the biplot basis in Fig. 10(d2) shows the inter-spin correlation captured by the biplot basis. The biplot basis capture the (anti)correlation of spins 1, 2, 5, 8, and 9 [yellow and blue arrows in Fig. 10(a2) top right] which corresponds to the left ring of the interaction network in Fig. 10(d2). The biplot basis also capture the (anti)correlation of spins 3, 4, 7, 10, and 11 [green and purple arrows in Fig. 10(a2) top right] which corresponds to the right ring of the interaction network in Fig. 10(d2). The hypercubic energy landscape using PCA provides insight into the dynamics through the correlation of spins.
@@ VII.5 The hypercubic energy landscape of the three-ring system
Unlike the one-ring and two-ring systems, the structure of pathways is more complex, but we can still infer the state transition from the projection. For example, the pathway from the ground state $\begin{bmatrix}\circ&\bullet&\circ&\bullet&\bullet&\circ&\circ&\circ&\circ&\bullet&\bullet&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ (around $\begin{bmatrix}-3&0.9\end{bmatrix}^{\top}$) to $\begin{bmatrix}\circ&\bullet&\bullet&\circ&\bullet&\circ&\bullet&\circ&\circ&\bullet&\circ&\bullet&\circ&\circ&\bullet\end{bmatrix}^{\top}$ (around $\begin{bmatrix}-1.5&3.2\end{bmatrix}^{\top}$) is a transition with the correlated spin flips of spins located in the lower right hexagon (spins 3, 4, 7, and 11). Pathways parallel to this pathway involve the same spin flips. From the biplot basis on the top right of Fig. 10(a3), one can generally know that by following the angle of each biplot basis, the transitions involve spins on the top hexagon (spins 12, 13, 14, and 15) being flipped: the pathways are lines from the left top to the right bottom, which are parallel to the biplot basis of spins 12, 13, 14, and 15. Likewise, horizontal pathways are used when spins in the lower left hexagon (spins 1, 2, 5, and 8) are flipped, and vertical pathways are followed when spins in the inverted triangle at the center (spins 6, 9, and 10) are flipped. Similar to the one-ring and two-ring systems, the pathway is a collective flip of spins reflecting the interaction network. Notice that this projection of the hypercubic energy landscape [Fig. 10(a3)] is a partial Hamming projection: the PC1 score is equivalent to the Hamming distance from the ground state on the left side (around $\begin{bmatrix}-3&0\end{bmatrix}^{\top}$) or the right side (around $\begin{bmatrix}3&0\end{bmatrix}^{\top}$) but without considering spins 6, 9, and 10.
@@
To understand the relation between PC loading and the interaction network, in particular, the relation between PC loading and the frustration of the system, we color the loading of PC1 and PC2 in Figs. 10(b3) and 10(c3), following the angle of the biplot basis in Fig. 10(a3). We confirmed that the element-wise sign of PC1 [Fig. 10(b3)] corresponds to the ground state group containing six ground states, but the states of spins 6, 9, and 10 are not determined. These spins involve a frustrated cycle of the interaction network. PC2 loading in Fig. 10(c3) shows that spins 1, 2, 5, and 8 do not contribute, and those spins belong to the lower left ring. Moreover, PC2 loading of spins 6, 9, and 10 has slightly higher or lower magnitudes than the others, indicating the uncommon contribution of spins 6, 9, and 10, suggesting the frustration-related correlation.
@@
These insights are consistent with the interaction network with colored nodes in Fig. 10(d3): the PC loading captures the correlation arising from the interaction network. Nevertheless, spins 6 and 9 have exactly the same color: their biplot basis are the same, but it can be another combination of spins such as 6 and 10. This breaks the symmetry of the system—PC2 ignores the lower left hexagon of the interaction network even though the system has the same other structure—and we examine how the PC loading reflects the symmetry of the system.
VII.6 Fraction of explained variance
We investigate the fraction of explained variance by the PCs of the hypercubic energy landscape in Fig. 11 to see how the frustration of the system influences the PCs. For the one-ring system [Fig. 11(a1)], PC1 dominates the explained variance by the PCs, which is consistent with PC1 being proportional to the ground state. Turning to the two-ring and three-ring systems [Figs. 11(a2) and 11(a3)], the first PCs explain less than 90% of the variance. This decrease in the explained variance by the first component arises from the degeneracy of the ground states. PC2 explains the variance of degenerate ground states. For the three-ring system [Fig. 11(a3)], the first three PCs are required to explain more than 80% of the variance, and the second and third fractions of explained variance are the same, indicating that PC2 and PC3 share the regularity: the symmetry of the system.